Compute a 95% confidence interval for the population mean, based on the numbers 1, 2, 3, 4, 5, 6, 20. Change the number 20 to 7 and recalculate the confidence interval. Using these
results, describe the effect of an outlier (i.e. extreme value) on confidence interval.
1
Expert's answer
2020-11-26T19:16:09-0500

"1.\ s=\sqrt{\frac{\sum(x_i-\overline{x})^2}{n-1}}\text{ --- sample variance}.\\
\text{Using Excel (STDEV.S function) we can compute that } s\approx 6.4660.\\
\overline{x}=\frac{41}{7}\\
\text{Critical value } t_{0.05}\approx 2.4469\text{ (we use t-distribution table)}\\
\text{We have:}\\
(41/7-(2.4469)(\frac{6.4660}{\sqrt{7}}),41/7+(2.4469)(\frac{6.4660}{\sqrt{7}}))\\
(-0.1229,11.8372)\\
2.\ s=\sqrt{\frac{\sum(x_i-\overline{x})^2}{n-1}}\text{ --- sample variance}.\\
\text{Using Excel (STDEV.S function) we can compute that } s\approx 2.1602.\\
\overline{x}=4\\
\text{Critical value } t_{0.05}\approx 2.4469\text{ (we use t-distribution table)}\\
\text{We have:}\\
(4-(2.4469)(\frac{2.1602}{\sqrt{7}}),4+(2.4469)(\frac{2.1602}{\sqrt{7}}))\\
(2.0022,5.9978)\\
3.\ \text{The confidence interval is wider in the first case}\\
\text{because of the outlier 20}."